OpenStudy (anonymous):

Convert to a decimal: 17/25

OpenStudy (anonymous):

just divide them

OpenStudy (anonymous):

It's 0.68. :)

OpenStudy (anonymous):

17/25 =(4×17)4×25 =68/100 =0.68

OpenStudy (anonymous):

thanks to both of you

OpenStudy (anonymous):

Yup np

OpenStudy (anonymous):

You are welcome. :)

OpenStudy (cwrw238):